OpenStudy (anonymous):

If two standard number cubes are rolled 100 times, predict the number of times that a sum of 7 or 11 will be rolled

OpenStudy (unklerhaukus):

with two standard diec the chances of a seven is 6 in 36 or 1/6 in 100 roles you can expect 100/6 = 16.66... ~ 17 sevens the chances of an eleven is 2 in 36 or 1/18 in 100 roles you can expect 100/18 = 5.55... ~ 6 elevens

OpenStudy (unklerhaukus):

i pulled 6 and 2 from a similar table

Directrix (directrix):

Looking at the uploaded image of possible outcomes of tossing a pair of dice once, notice that the sums of 7 are aligned on a right to left diagonal. There are 6 ways to get a sum of 7. The sums of 11 are on a mini-diagonal, right to left. There are two ways to get 11. That gives a total of 8 ways per toss to get a sum of 11 or a sum of 7. The associated probabilty is 8/36 for one roll. For 100 rolls, 100* (8/36) = 22.22 approx. Based on that, I would expect 22 sums of 7 or 11 in 100 rolls of a pair of dice.